what is cost of metal in a single can if radius is 1.30 inches height is 3.95 inches and cost is 0.0040.
Found 2 solutions by Alan3354, Boreal:
Answer by Alan3354(69443)

what is cost of metal in a single can if radius is 1.30 inches height is 3.95 inches and cost is 0.0040.
----------
Cost is 0.0040.
That's stated above.

Answer by Boreal(15235)

Area is 2*pi*r^2+2pi*rh=42.88 in^2*(0.0040 per sq in, I assume)=0.1715, cost units not known.
Check with 2pi*r*h+2 pi*r^2=2pi*r(h+r)=42.88